[figure: 9.7 cm]

Synthesis, Structure and Magnetic Properties of Pb_{2}VO(PO_{4})_{2}: The

second example for a frustrated S=1/2 square lattice Antiferromagnet

E. E. Kaul^{a}, C. Geibel^{a}, A. V. Mironov^{b}, R. V. Shpanchenko^{b} and E. V.

Antipov^{b} aMax-Planck Institut für Chemische Physik fester Stoffe,Dresden, Germany bMoscow State University, Dep. of Chemistry, Moscow, Russia

Among low dimensional spin systems, frustrated S=1/2 square lattice systems have attracted strong theoretical interest, since an interesting phase diagram and unusual ground states are expected depending on the strength of the diagonal frustrating exchange J_{2}. However, experimental studies were very limited since the only compound where such spin system is realized (Li_{2}VOA^{+4}O_{4} (A=Si, Ge)) was discovered only 2 years ago. We have discovered a new compound, Pb_{2}VO(PO_{4})_{2} where the structural data as well as susceptibility and specific heat measurements suggest that such a frustrated S=1/2 lattice is also realized. Pb_{2}VO(PO_{4})_{2} has a structure formed by layers composed by S=1/2 V^{+4} in pyramidal coordination connected by PO_{4} tetrahedra similar to that found in Li_{2}VOA^{+4}O_{4} (A=Si, Ge) but modulated along b. These layers are well separated by layers of Pb atoms and isolated PO_{4} tetrahedra. Magnetic Susceptibility measurements shows a clear maximum at T_{Max} ≈ 8.7 K with a decrease towards lower temperatures. At T ≈ 3.6K we observe a very weak magnetic transition which becomes very pronounced and shifts to higher temperatures with applied magnetic field. This suggest that Pb_{2}VO(PO_{4})_{2} is very close to a quantum critical point.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : E. E. Kaul

Address for correspondence : Nöthnitzer Str. 40 01187 Dresden Germany

Email address : kaul@cpfs.mpg.de